The x intercept of the graph of 2x+3y=5 is

Please show all work
Thank You

Q23

Answer by chitra(359)

The x intercept of the graph of 2x + 3y = 5
To find the x intercept of any equation plug in y = 0 in any equation.
So when y = 0 in the given equation, gives us
2x + 3(0) = 5
2x + 0 = 5
2x = 5
x = 5/2
Thus, the solution